OpenStudy (anonymous):

Let vector AB=(6,-2,3)and vector AC= (-2,-3,2). How do I find BÂC?

OpenStudy (anonymous):

AB+BC+CA=0

OpenStudy (anonymous):

do you understand why?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

draw a triangle. Label the three vertices A B and C. See that if you move from A to B, then from B to C, then from C to A, you end up right where you started: at A.

OpenStudy (anonymous):

 .ok I see that.

OpenStudy (anonymous):

so the total vector change (going from A to B, from B to C, and from C back to A) AB+BC+CA=0 (you end exactly where you started)

OpenStudy (anonymous):

OH. OK I understand,but how do I substitute the coordinates into that equation?

OpenStudy (anonymous):

you have AB and AC. Look back at your triangle. If you move from A to C and then from C back to A, you end up right where you started, right? so AC + CA = 0, so AC = -CA

OpenStudy (anonymous):

so then you have AB and CA. Substitute them into the equation and you can find BC by algebra.

OpenStudy (anonymous):

hold on, I just realize I wrote the problem wrong. I'm trying to find angle "a" not vector BC. I'm sorry

OpenStudy (anonymous):

do you know how to take the dot product of two vectors?

OpenStudy (anonymous):

no, my teacher have not taught my class how to multiply vectors. we just started. *BÂC

OpenStudy (anonymous):

the easy way to find the angle between the two sides is with the dot product. The more complicated way uses the Law of Cosines.

OpenStudy (anonymous):

I know the Law of cosine. we just quizzed on that and law of sine.

OpenStudy (anonymous):

so to find the angle with law of cosines, find BC, find the magnitudes of all three vectors, and set up the triangle. The side lengths are the magnitudes. You can then use the law of cosines to find the angle

OpenStudy (anonymous):



OpenStudy (anonymous):

How would you find the magnitudes?

OpenStudy (anonymous):

it's kind of like using the pythagorean theorem twice. Once to find out the magnitude of just the x coordinate and the y coordinate, then again to find out the magnitude of the magnitude you found before and the z coordinate. \[distance=\sqrt{\sqrt{x^2+y^2}^2+z^2}\]simplifies to \[distance=\sqrt{x^2+y^2+z^2}\]

OpenStudy (anonymous):

(sorry it took so long)

OpenStudy (anonymous):

no I'm sorry I keeping looking at the wrong problem. I am trying to find vector BC. I'm so sorry I making you all this work. I'm such a mess.

OpenStudy (anonymous):

it's fine

OpenStudy (anonymous):

I would not be surprise if you stop helping me . I understand

OpenStudy (anonymous):

you already know how to do this, I showed you up above.

OpenStudy (anonymous):

AB+BC+CA=0 just use algebra

OpenStudy (anonymous):

I meant to write how do I substitute AB and AC into the equation?

OpenStudy (anonymous):

CA is -AC, so CA is (2,3,-2)

OpenStudy (anonymous):

has your teacher taught you how to add vectors?

OpenStudy (anonymous):

I understand that. When I substitution the coordinates, how would I write it. That is the part where I am confused at. She just began teacher us vectors, like a seven min intro. So, my class is basically got nowhere.

OpenStudy (anonymous):

she like and always makes us figure it out ourselves.

OpenStudy (anonymous):

I write vectors either as \[\left(\begin{matrix}x \\ y \\ z \end{matrix}\right)\] or as (x,y,z)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the first form is better if you want other people to understand it, but the second form is faster and takes less paper

OpenStudy (anonymous):

well, which way do u want to do it?

OpenStudy (anonymous):

it's your math problem, you choose how you write it but your teacher would probably be happier if you wrote it vertically so you have \[\left(\begin{matrix}6 \\ -2 \\ 3\end{matrix}\right) + \left(\begin{matrix}2 \\ 3 \\ -2\end{matrix}\right)=-BC\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so, does BC equal (-8,-1,-1)?

OpenStudy (anonymous):

looks like it does

OpenStudy (anonymous):

yah. thankyou so much!